OpenStudy (abbles):

Find an equation of the tangent to the curve at the given point by two methods: (a) without eliminating the parameter and b) by first elminating the parameter

OpenStudy (abbles):

\[x = 1 + lnt, y = t^2 + 2\]

OpenStudy (abbles):

At (1, 3)

zepdrix (zepdrix):

For part a,\[\large\rm \frac{dy}{dx}=\frac{\dot y}{\dot x}\]

OpenStudy (abbles):

I know how to do this problem given a value for t, but not given the point...

OpenStudy (abbles):

For dy/dx I got 2t^2

OpenStudy (abbles):

What are the dots on top of the y and x?

zepdrix (zepdrix):

It's just a different notation used for `time derivatives`. \(\large\rm \frac{dy}{dt}=\dot y=y'\)

OpenStudy (abbles):

Ah interesting. Okay

zepdrix (zepdrix):

\[\large\rm \color{orangered}{x}=1+\ln t\qquad\qquad\qquad \color{orangered}{y}=t^2+2\]We have values for x and y at least,\[\large\rm \color{orangered}{1}=1+\ln t\qquad\qquad\qquad \color{orangered}{3}=t^2+2\]

zepdrix (zepdrix):

This should allow you to find values for t.

zepdrix (zepdrix):

ln(t)=0 has no solution. Are you sure you have the equations written down properly? Hmm

OpenStudy (abbles):

Hmm yes the equations are correct. The book gives the answer as: y = 2x + 1

zepdrix (zepdrix):

Did I say ln(t)=0 has no solution? Oh boy my brain is failing today :) lol. I meant ln(0) has no solution. But obviously ln(t)=0 has the solution of t=1, right?

zepdrix (zepdrix):

You have to make sure that value satisfies both equations though.

zepdrix (zepdrix):

Your derivative looks correct.\[\large\rm \frac{dy}{dx}=2t^2\]

zepdrix (zepdrix):

Evaluating this derivative at x=1, y=3 and t=1,\[\large\rm \frac{dy}{dx}=2\]This gives us the `slope` of the tangent line, ya?\[\large\rm y-y_o=m(x-x_o)\]\[\large\rm y-y_o=2(x-x_o)\]

OpenStudy (abbles):

Where did you get the values x=1, y=3 and t=1 from?

zepdrix (zepdrix):

(x,y) = (1,3) corresponds to some `specific t value`.

zepdrix (zepdrix):

x and y take on these values at some `specific time`.

zepdrix (zepdrix):

By plugging x=1 and y=3 into our equations, we are able to solve for t.

zepdrix (zepdrix):

\[\large\rm \color{orangered}{x}=1+\ln t\qquad\qquad\qquad \color{orangered}{y}=t^2+2\]\[\large\rm \color{orangered}{1}=1+\ln t\qquad\qquad\qquad \color{orangered}{3}=t^2+2\]

zepdrix (zepdrix):

\[\large\rm 3=t^2+2\]Subtracting 2,\[\large\rm 1=t^2\]Square root,\[\large\rm t=\pm1\]So t could potentially be 1 or negative. We need to find the value that satisfies BOTH equations though.\[\large\rm 1=1+\ln t\]\[\large\rm 0=\ln t\]t=-1 will NOT satisfy this equation. So t=1 is the specific time that we were looking for.

OpenStudy (abbles):

Gotcha! Now all that's left to do is find the slope at t = 1, which is 2. y - 3 = 2(x-1) y = 2x +1 Yep :) thanks zep!

OpenStudy (abbles):

What about for part b though?

zepdrix (zepdrix):

\[\large\rm x=1+\ln t\qquad\qquad\qquad y=t^2+2\]Hmm.. Let's try solving for t in the first equation.\[\large\rm x-1=\ln t\]\[\large\rm e^{x-1}=t\]And I guess we can replace the t in our second equation with this, ya?

OpenStudy (abbles):

Yes that sounds right

OpenStudy (abbles):

\[y = e ^{2x-2}+2\]

OpenStudy (abbles):

Like this right?

zepdrix (zepdrix):

Umm yes. So doing it this way, you've completely eliminated the parameter. No need for your fancy parameter derivative.. thing.

OpenStudy (abbles):

So would I take the derivative the regular way?

zepdrix (zepdrix):

Yes

OpenStudy (abbles):

Does this look right? \[(4x-4)e ^{2x-2}\]

zepdrix (zepdrix):

\[\large\rm y'=e^{2x-2}(2x-2)'\]

OpenStudy (abbles):

Wouldn't it need to be multiplied by the inside function? So by 2?

OpenStudy (abbles):

*The derivative of the inside function I mean

zepdrix (zepdrix):

\[\large\rm y'=e^{2x-2}(2)\]Ya that looks better.

OpenStudy (abbles):

... but the top still needs to come down right? Like \[y' = (2x-2)e ^{2x-2}*2\]

OpenStudy (abbles):

Or am I wrong?

zepdrix (zepdrix):

No, the top doesn't come down. The derivative of the top comes down only.

zepdrix (zepdrix):

Gotta master that chain rule girly!! :)

OpenStudy (abbles):

Mmmmm yeah, the e^x and e^a stuff always trips me up. So \[2e ^{2x-2}\] Then what?

zepdrix (zepdrix):

\(\large\rm y'(x)=2e^{2x-2}\) Then evaluate your derivative function at x=1, y=3.

OpenStudy (abbles):

By plugging in x =1?

OpenStudy (abbles):

It would equal 2. Then the same steps as last time to get the equation. Alrighty :) Zepman to the rescue.

zepdrix (zepdrix):

yayyy team

OpenStudy (abbles):

:D